OpenStudy (anonymous):

A sample of 143 golfers showed that their average score on a particular golf course was 80.35 with a standard deviation of 3.33. Answer each of the following (show all work and state the final answer to at least two decimal places.): (A) Find the 95% confidence interval of the mean score for all 143 golfers. (B) Find the 95% confidence interval of the mean score for all golfers if this is a sample of 90 golfers instead of a sample of 143. (C) Which confidence interval is smaller and why? (Points : 6)

OpenStudy (kropot72):

The formula to find ther answer to (A) is: \[xbar-1.96\frac{\sigma}{\sqrt{n}}< \mu < xbar+\frac{\sigma}{\sqrt{n}}\] You just need to substitute. Can you do that?